OpenStudy (anonymous):

Question:

OpenStudy (anonymous):

Using: \[\sum_{r=0}^{n}\left(\begin{matrix}n \\ r\end{matrix}\right)rx^{r}=nx(1+x)^{n-1}\] Show what: \[\sum_{r=1}^{n}\left(\begin{matrix}n \\ r\end{matrix}\right)r^{2}=n(n+1)2^{n-2}\]

OpenStudy (anonymous):

Show that*

OpenStudy (anonymous):

@ParthKohli: \[(1+x)^{n} = \sum_{r=0}^{n}\left(\begin{matrix}n \\ r\end{matrix}\right)x^{r}\]

OpenStudy (anonymous):

Want to learn the proofs Parth?

Parth (parthkohli):

Yes sir!

OpenStudy (anonymous):

Okay, do you know this proof: \[\left(\begin{matrix}n \\ r\end{matrix}\right)+ \left(\begin{matrix}n \\ r+1\end{matrix}\right)= \left(\begin{matrix}n+1 \\ r+1\end{matrix}\right)\]?

Parth (parthkohli):

I guess it goes something like: \( \color{Black}{\Rightarrow \Large {n! \over (n - r)!r!} + {n! \over (n - r - 1)!(r -1)!}}\)

OpenStudy (anonymous):

Bingo! To simplify it; you must know these: \((r+1)! = (r+1)*r!\) \((n-r)!= (n-r)(n-r-1)!\) Are you able to proceed further?

OpenStudy (anonymous):

Remember; to find the common factor!

Parth (parthkohli):

I am afk. Can I talk to you later?

OpenStudy (anonymous):

Hang on..it's not right.. This is the correct one. \[\color{Black}{\Rightarrow \Large {n! \over (n - r)!r!} + {n! \over (n - r - 1)!(r +1)!}}\]

OpenStudy (anonymous):

Okay; If I'm still here lol